OpenStudy (anonymous):

if 5 out of 25 people like choclate ice cream, what percent of the people do not like choclate ice cream. A.20% B.5% C.80% D.30%

OpenStudy (anonymous):

IWILL GIVE A MEADEL

OpenStudy (anonymous):

B, I think. :) I'm pretty sure actually.

OpenStudy (anonymous):

IS IT B

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

It's A, sorry. I read the question wrong. I though it meant people who like chocolate ice cream.

OpenStudy (anonymous):

why how do u know

OpenStudy (anonymous):

It is A!!!!!!

OpenStudy (anonymous):

20% of 25 = 5

OpenStudy (anonymous):

So do you get it?

OpenStudy (i<3u):

25 - 5 = ? @oregon234

OpenStudy (anonymous):

are u guys sure um this is like really important

OpenStudy (anonymous):

Yh it is correct. A is the answer

OpenStudy (anonymous):

@oregon234 A is it. I did not mean B. I was wrong at first, I read the question wrong. It is A!

OpenStudy (anonymous):

@CrazyCountryGirl Can you help me with my question?

OpenStudy (anonymous):

Yeah, just tag me. :)

OpenStudy (anonymous):

I did :3

OpenStudy (anonymous):

5 out of 25 people like ice cream, then 20 out of 25 people don't like it. 20/25 = 4/5 = .8 = 80% So, 80% of the people DON'T like it. 20% of the people DO like it.